# Convergence and Generalization of Anti-Regularization for Parametric Models

### v1

@@ Abstract
We propose Anti-regularization (AR), which adds a sign-reversed reward term to the loss to intentionally increase model expressivity in the small-sample regime, and then attenuates this intervention with a power-law decay as the sample size grows. We formalize spectral safety and trust-region conditions, and design a lightweight stability safeguard that combines a projection operator with gradient clipping, ensuring stable intervention under stated assumptions. Our analysis spans linear smoothers and the Neural Tangent Kernel (NTK) regime, providing practical guidance on selecting the decay exponent by balancing empirical risk against variance. Empirically, AR reduces underfitting while preserving generalization and improving calibration in both regression and classification. Ablation studies confirm that the decay schedule and the stability safeguard are critical to preventing overfitting and numerical instability. We further examine a degrees-of-freedom targeting schedule that keeps per-sample complexity approximately constant. AR is simple to implement and reproducible, integrating cleanly into standard empirical risk minimization pipelines. It enables robust learning in data- and resource-constrained settings by intervening only when beneficial and fading away when unnecessary.
@@ 1.1 Motivation
While the adoption of artificial intelligence (AI) in industry has become widespread, small- and medium-sized enterprises (SMEs) and startups often begin with limited samples due to constraints in computing resources and data infrastructure. These disparities and practical limitations have been repeatedly highlighted in international comparisons. According to OECD reports [1, 2], there are pronounced gaps in AI adoption and capabilities across industries, regions, and firm sizes. In particular, SMEs suffer from structural bottlenecks in data acquisition, quality control, and operational deployment in early-stage AI projects, due to limited computing resources, poor data governance, and insufficient talent capacity. These realities suggest the need for a training procedure that delivers immediately usable performance while avoiding underfitting, especially under the assumption that real-world applications often begin with small samples.
@@
In data-scarce environments, acquiring additional samples can be prohibitively costly and time-consuming. Privacy regulations and industry norms also hinder data sharing and integration. The European Parliamentary Research Service [3] notes that GDPR imposes constraints on AI training data through the principles of data minimization and purpose limitation. In practice, small-sample learning scenarios are common. Tian et al. [4] systematize methodologies for incremental data settings, and Tao et al. [5] emphasize the intrinsic scarcity of samples in rare-event tasks such as industrial defect detection. While data augmentation is an important tool, it is not a panacea. Wang et al. [6] summarize both the benefits and limitations of augmentation, while Geirhos et al. [7] and Brown et al. [8] show how spurious correlations and shortcut learning can hinder generalization in real-world deployments.
@@
In such early stages—where data is limited and risks of noise and bias coexist—traditional regularization, which suppresses model complexity, may exacerbate underfitting. Hastie et al. [9] review the bias–variance tradeoff and the impact of regularization in small-sample regimes, and Belkin et al. [10] demonstrate that expanding expressivity or relaxing constraints can, under certain conditions, improve generalization. In practice, the initial model performance—what is immediately available—often determines subsequent investment decisions and project continuation. Paleyes et al. [11] and Sculley et al. [12] highlight the criticality of early-stage performance validation, considering deployment constraints, technical debt, and cost structures.
@@
This study aims to address such practical demands by proposing a training strategy that secures usable performance even in small-sample regimes, while enabling a stable transition to the final model without changing the architecture as data accumulates. Unlike continual learning, we perform retraining at each stage, adjusting the learning strategy according to the sample size to manage the overall training trajectory in a continuous manner. The conceptual distinction from continual learning can be referenced in De Lange et al. [13]. Since reversing the sign of a regularization term carries risks of divergence, we design safety mechanisms to ensure stable reward operation: spectral safety conditions for regression and trust-region constraints with bounded margin rewards for classification. Schulman et al. [14] and Cortes and Vapnik [15] offer the standard frameworks for trust-region constraints and margin-based design, respectively. The reward strength $\lambda$ follows a decay schedule that gradually diminishes with the sample size $n$, so that as data accumulates, the effect of the reward naturally vanishes, ultimately converging to training without a regularization term.
@@ 1.2 Contributions
This study proposes Anti-regularization (AR), which adds a reward term to the loss function, to achieve immediately usable performance even in small-data and resource-constrained environments. The proposed method selectively enhances expressivity in the early stages to mitigate underfitting and gradually attenuates the reward strength as the sample size increases, enabling a smooth transition to a final model without regularization. The main contributions are as follows:
@@
  • Theoretical formalization and safety conditions of Anti-regularization: We formalize Anti-regularization, which introduces a reward term with a reversed sign into the loss, and establish design principles that prevent divergence in both regression and classification tasks. We also discuss its extensibility to shallow neural networks and the Neural Tangent Kernel regime.
@@
  • Sample-size–dependent schedules and DoF targeting: We propose a simple schedule that gradually reduces the reward strength with increasing sample size, and present an alternative degrees of freedom–based rule that maintains constant per-sample complexity.
@@
  • Practical diagnostics for alleviating underfitting: We outline a procedure to determine when to activate or deactivate Anti-regularization based on the joint trends of training error and model complexity, and provide monitoring metrics to verify its contribution to actual performance improvement.
@@
  • Safe training protocol for classification tasks: To avoid instability due to the absence of lower bounds, we construct a robust training procedure that combines trust-region constraints, bounded margin rewards, logits, and gradient clipping.
@@
  • Lightweight and compatible implementation: Targeting linear models and shallow multilayer perceptrons, we propose a practical protocol that boosts initial performance by adjusting only the loss function without modifying model architecture or hardware, then naturally transitions to a no-regularization model as data accumulates.
@@
  • Systematic experiments for theoretical validation: Focusing on tabular regression and image classification, we evaluate the validity, safety, and sensitivity of the proposed method and empirically verify its effectiveness through comparisons with alternative strategies.
@@
  • Practical alternative for small organizations: By securing reasonable initial performance without requiring high-performance computing or large-scale datasets, and by providing a transition path within a fixed model structure as data grows, we lower the barriers to AI adoption for SMEs and startups.
1.3 Paradigm Shifts in Regularization
@@
Zhang et al. [16] challenged classical intuitions on generalization by showing that large neural networks can perfectly fit randomly labeled data without explicit regularization, while still achieving non-trivial generalization. Soudry et al. [17] proved that gradient descent implicitly converges to the maximum-margin separator in linearly separable settings, revealing that the learning algorithm itself can induce a strong implicit regularization effect. Bartlett et al. [18] formalized the phenomenon of benign overfitting, where interpolating solutions in high-dimensional linear regression can still achieve test risk close to the noise level.
@@
Belkin et al. [10] introduced the double descent curve, showing that as model capacity increases, test risk can decrease again after passing through the traditional U-shaped bias–variance tradeoff curve, highlighting a structural shift in generalization behavior around the interpolation threshold. Loog et al. [19] revisited literature from statistics and pattern recognition in the 1980s–1990s, showing that double descent phenomena had been previously observed and that modern results lie on a historical continuum rather than being an abrupt shift. Nakkiran et al. [20] systematically studied deep double descent along axes such as model width, training epochs, and data size. Adlam et al. [21] emphasized that a fine-grained bias–variance decomposition is key to understanding the underlying mechanism of double descent.
@@
Hastie et al. [22] provided a rigorous risk analysis of ridgeless minimum-norm interpolators, showing that such solutions can generalize surprisingly well in overparameterized regimes. Belkin et al. [23] used two models to demonstrate how double descent arises under weak feature conditions. Muthukumar et al. [24] derived the harmless interpolation condition, showing that interpolation may not harm generalization when noise is spread across many less relevant directions.
@@
Geiger et al. [25] proposed a jamming transition between under- and overparameterized regimes, explaining the spike in generalization error near the interpolation threshold and its subsequent decline using insights from loss landscape geometry. Spigler et al. [26] independently observed and analyzed similar transitions and generalization patterns, reinforcing their reproducibility. Advani et al. [27] used high-dimensional mean-field theory to quantify how the data-to-model ratio and signal-to-noise ratio mediate such transitions. Meanwhile, Nakkiran et al. [28] theoretically and empirically showed that the choice and strength of optimal (especially $\ell_{2}$) regularization can mitigate double descent, suggesting that regularization should be contextually designed rather than uniformly strong.
@@ 1.4 Anti-regularization and Reward-Based Objectives
Unlike traditional regularization, which suppresses model complexity by adding penalties to the loss function, reward-based objectives—adding terms that promote specific behaviors or properties—have been explored across various contexts. Pereyra et al. [29] introduced a confidence penalty to reduce overconfident predictions, and Szegedy et al. [30] used label smoothing to soften the target distribution and suppress excessive certainty. Müller et al. [31] systematically analyzed the effects of smoothing, highlighting its benefits in terms of generalization and uncertainty calibration. Kim et al. [32] proposed negative learning, which uses incorrect labels as learning signals to enhance robustness in noisy-label settings, while Ishida et al. [33] restructured the objective to rely on complementary labels instead of correct ones.
@@
In sequence learning, efforts have been made to directly optimize task-specific rewards. Ranzato et al. [34] proposed learning sequence-level rewards using REINFORCE, and Norouzi et al. [35] generalized maximum likelihood training into reward-aligned learning via reward-proportional distributions. In language model fine-tuning guided by human preferences, Christiano et al. [36] and Ziegler et al. [37] proposed policy improvement via reward models trained on human feedback, and Ouyang et al. [38] demonstrated their effectiveness in large-scale instruction-following. More recently, Rafailov et al. [39] proposed a supervised alternative that implicitly maximizes rewards from preference data without explicit reinforcement learning.
@@
In representation learning, contrastive learning adopts a reward structure that promotes similarity among positive pairs and separation among negative pairs. van den Oord et al. [40] proposed a contrastive loss that maximizes a lower bound on mutual information, encouraging high-information representations. On the other hand, adversarial objectives designed to eliminate unwanted factors are closer to anti-reward mechanisms. Ganin et al. [41] used gradient reversal to make the domain classification loss act in reverse, preventing domain information from being encoded. Zhang et al. [42] trained with an adversary to remove bias signals by targeting sensitive attribute predictors. In reinforcement learning, Haarnoja et al. [43] incorporated entropy rewards into the reward function, formalizing the maximum entropy objective to enhance exploration and robustness.
1.5 Sample-Size–Dependent Regularization and Degrees of Freedom Control
@@
The classical idea of adjusting regularization strength based on sample size and targeting a desired level of effective degrees of freedom has been established in information criteria, risk estimation, and cross-validation theory. Akaike [44] penalized model complexity (number of effective parameters) as a constant term in the AIC, correcting bias in generalization error. Mallows [45] introduced the $C_{p}$ criterion as an unbiased estimate of prediction error. Stein [46] enabled data-adaptive shrinkage via Stein’s Unbiased Risk Estimate (SURE), and Craven et al. [47] formalized generalized cross-validation (GCV) for choosing smoothing parameters in a data-dependent manner. Wahba [48] established the equivalence between degrees of freedom and smoothness control in spline regularization. Ye [49] defined generalized degrees of freedom (GDF) as the sum of sensitivities of fitted values to observations and showed that, in linear smoothers, it matches the trace of the hat matrix—providing a theoretical foundation for data-driven complexity penalties. Hastie et al. [9] unified these criteria from a common perspective and discussed practical rules for gradually relaxing penalties as the sample size increases to avoid underfitting and minimize prediction risk.
@@
In sparsity-inducing regularization, the interplay between degrees of freedom and penalty strength has also been rigorously analyzed. Efron et al. [50] described the piecewise linear structure of the Lasso path and how degrees of freedom vary along it. Zou et al. [51] showed that the degrees of freedom of the Lasso can be approximated by the number of active variables under general conditions. Tibshirani et al. [52] characterized the degrees of freedom of Lasso-type problems under convexity and regularity assumptions, providing quantitative guidance for scaling regularization with respect to sample size and signal strength. Donoho et al. [53] demonstrated that SURE-based threshold selection enables data-dependent shrinkage that minimizes risk, and Stone [54] theoretically justified cross-validation as a mechanism for balancing underfitting and overfitting as sample size increases.
@@
At the level of training procedures, early stopping has been shown to act as a data-dependent form of regularization. Prechelt [55] empirically demonstrated that using validation error to define stopping criteria suppresses overfitting by providing a regularization effect. Yao et al. [56] showed that early stopping in gradient descent can be equivalent to Tikhonov regularization under certain conditions. Raskutti et al. [57] derived generalization bounds in nonparametric regression when the number of iterations is chosen based on sample size, supporting the conclusion that longer training or weaker explicit regularization becomes favorable as data grows.
@@
In stochastic optimization, the noise scale induced by batch size and learning rate has been increasingly recognized as an implicit regularization mechanism, which must be co-scheduled with sample size. Mandt et al. [58] approximated SGD as a stochastic differential equation and showed that noise temperature plays a regularization role by mediating between prior and posterior. Smith et al. [59] explained optimal generalization using a scaling law between batch size, learning rate, and dataset size. Jastrzębski et al. [60] reported how combinations of data size, learning rate, and batch size influence flat minima and generalization. Keskar et al. [61] pointed out that large batches may weaken implicit regularization, leading to generalization gaps. Loshchilov et al. [62] decoupled weight decay from optimization steps to enable better scheduling, and Smith [63] proposed cyclical learning rates to dynamically control the strength of implicit regularization in practice.
1.6 Classification Safety and Training Stability

### v2

@@ Abstract
Anti-regularization introduces a reward term with a reversed sign into the loss function, deliberately amplifying model expressivity in small-sample regimes while ensuring that the intervention gradually vanishes as the sample size grows through a power-law decay schedule. We formalize spectral safety conditions and trust-region constraints, and we design a lightweight safeguard that combines a projection operator with gradient clipping to guarantee stable intervention. Theoretical analysis extends to linear smoothers and the Neural Tangent Kernel regime, providing practical guidance on the choice of decay exponents through the balance between empirical risk and variance. Empirical results show that Anti-regularization mitigates underfitting in both regression and classification while preserving generalization and improving calibration. Ablation studies confirm that the decay schedule and safeguards are essential to avoiding overfitting and instability. As an alternative, we also propose a degrees-of-freedom targeting schedule that maintains constant per-sample complexity. Anti-regularization constitutes a simple and reproducible procedure that integrates seamlessly into standard empirical risk minimization pipelines, enabling robust learning under limited data and resource constraints by intervening only when necessary and vanishing otherwise.
@@ 1.1 Motivation
The adoption of artificial intelligence (AI) has become widespread in industrial practice, yet small and medium-sized enterprises (SMEs) and startups often begin with limited sample sizes due to constraints in computing resources and data infrastructure. Such disparities and practical limitations are repeatedly observed in international comparisons. According to OECD reports [1, 2], significant divides in AI adoption and capability exist across industries, regions, and firm sizes. In particular, SMEs face structural bottlenecks in data acquisition, quality assurance, and operational deployment stages of early AI projects due to limited computing and data resources, weak data governance, and insufficient expertise. These realities highlight the need for training procedures that can deliver immediately usable performance while avoiding underfitting, even when starting from very small samples.
@@
In severely data-constrained environments, acquiring additional samples entails substantial costs and delays, and data sharing or integration is further hindered by privacy regulations and industry norms. The European Parliamentary Research Service [3] notes that GDPR, through principles such as data minimization and purpose limitation, may restrict the use of training data for AI. In practice, small-sample learning scenarios are frequent. Tian et al. [4] systematize methodological approaches under incremental small-sample regimes, and Tao et al. [5] summarize how tasks such as industrial defect detection, which are driven by rare events, are inherently sample-limited. While data augmentation is an important tool, it is not a universal solution. Wang et al. [6] review both the benefits and limitations of augmentation, and Geirhos et al. [7] and Brown et al. [8] show that spurious correlations and shortcut learning can impair generalization in real-world deployments.
@@
At these early stages, where data scarcity coexists with the risks of noise and bias, traditional regularization that suppresses model complexity can exacerbate underfitting. Hastie et al. [9] provide an overview of the bias–variance trade-off and the role of regularization in small-sample regimes, while Belkin et al. [10] demonstrate that expanding expressivity or relaxing constraints can improve generalization under certain conditions. In practical terms, the initial performance of a model that is ready for immediate use often determines investment decisions and the continuation of projects. Paleyes et al. [11] and Sculley et al. [12] emphasize the importance of early validation through deployment constraints, operational requirements, technical debt, and cost structures.
@@
This study addresses these practical demands by pursuing a training strategy that secures usable performance even with small samples while enabling a stable transition to a final model as data accumulates, without altering the model architecture. Unlike continual learning, we assume retraining at each stage, but we adapt the training strategy according to the sample size, thereby managing the overall training trajectory in a continuous manner. For conceptual contrast, see De Lange et al. [13]. Since reversing the sign of a regularization term inherently entails a risk of divergence, we design safeguards so that the reward functions operate stably: in regression, by employing spectral safety conditions, and in classification, by employing trust-region constraints together with bounded margin rewards. Schulman et al. [14] and Cortes and Vapnik [15] provide the canonical frameworks for trust regions and margin bounds, respectively. The reward strength $\lambda$ follows a schedule that decays progressively with the sample size $n$, such that as data accumulates, the effect of the reward naturally vanishes, ultimately converging to training without any regularization term.
@@ 1.2 Contributions
This study proposes Anti-regularization (AR), which augments the loss function with a reward term, to achieve immediately usable performance under small-data and resource-constrained settings. The proposed approach selectively enhances expressivity during early stages with limited data to mitigate underfitting, while progressively decaying the reward strength as samples accumulate, thereby enabling a smooth transition to a final model trained without regularization. The main contributions are as follows:
@@
  • Theoretical formalization and safety conditions of Anti-regularization: We formalize Anti-regularization by introducing a reward term with reversed sign into the loss, and we establish design principles that prevent divergence in both regression and classification. We also discuss its extension to shallow neural networks and the Neural Tangent Kernel regime.
@@
  • Sample-size–dependent schedules and degrees-of-freedom targeting: We propose a simple schedule in which reward strength gradually decreases with the sample size, and as an alternative, we introduce a degrees-of-freedom targeting rule that maintains constant per-sample complexity.
@@
  • Practical diagnostics for mitigating underfitting: We outline a procedure to determine when to introduce and remove Anti-regularization by jointly monitoring training error and model complexity, and we provide metrics to assess whether it contributes to genuine performance improvements.
@@
  • Safe training protocol for classification tasks: To avoid instability caused by the absence of lower bounds, we design a stable training procedure that combines trust-region constraints, bounded margin rewards, logits, and gradient clipping.
@@
  • Lightweight and compatible implementation: Targeting linear models and shallow multilayer perceptrons, we propose a practical protocol that boosts early performance by modifying only a component of the loss, without changing architectures or hardware, and naturally transitions to unregularized models as data accumulates.
@@
  • Systematic experiments for theoretical validation: We evaluate validity, safety, and sensitivity on tabular regression and image classification tasks, and we verify the effectiveness of the proposed method through comparisons with alternative strategies.
@@
  • Practical alternative for small organizations: By securing a reasonable level of early performance without high-performance computing or large-scale data, and by providing a transition path within the same model architecture as data grows, we lower barriers to AI adoption for SMEs and startups.
2 Related Work
2.1 Shifts in the Regularization Paradigm
@@
Zhang et al. [16] demonstrated that large neural networks can perfectly fit even random labels without explicit regularization while still maintaining generalization performance, challenging classical intuitions about generalization. Soudry et al. [17] proved that gradient descent converges implicitly to the maximum-margin direction in linearly separable data, showing that the inductive bias of the learning algorithm itself exerts a powerful regularization effect. Bartlett et al. [18] formalized the phenomenon of benign overfitting, where in high-dimensional linear regression models, interpolation with zero training error can nonetheless yield test error close to the noise level.
@@
Belkin et al. [10] introduced the double descent curve, where test risk, instead of following the traditional U-shape, descends again as model capacity increases beyond the interpolation threshold, indicating a structural shift in generalization behavior before and after interpolation. Loog et al. [19] revisited statistical and pattern recognition literature from the 1980s and 1990s, systematizing earlier observations of phenomena resembling double descent and showing that modern findings stand in historical continuity rather than abrupt departure. Nakkiran et al. [20] provided a systematic exploration of deep double descent across multiple axes, including model size, training epochs, and dataset size. Adlam et al. [21] highlighted that a fine-grained bias–variance decomposition is essential to understanding the mechanisms underlying double descent.
@@
Hastie et al. [22] rigorously analyzed the risk of ridgeless minimum-norm interpolators, showing that in overparameterized regimes, minimum-norm solutions can unexpectedly achieve strong generalization. Belkin et al. [23] explained how double descent manifests in weak-feature settings through two canonical models. Muthukumar et al. [24] derived conditions for harmless interpolation, showing that interpolation need not be detrimental when noise is spread across many unimportant directions.
@@
From a loss landscape perspective, Geiger et al. [25] proposed the jamming transition between under- and overparameterization, physically explaining why generalization error peaks near the interpolation threshold and then decreases again in the overparameterized regime. Spigler et al. [26] independently observed and analyzed similar transitions and generalization patterns, thereby strengthening reproducibility. Advani et al. [27] used high-dimensional mean-field analysis to quantify how the data-to-model ratio and signal-to-noise ratio mediate these transitions. Meanwhile, Nakkiran et al. [28] showed both theoretically and empirically that the choice and strength of optimal (particularly $\ell_{2}$) regularization can mitigate double descent, suggesting that regularization must be designed adaptively rather than always enforced strongly.
@@ 2.2 Anti-regularization and Reward-based Objectives
Unlike traditional regularization, which suppresses model complexity by adding penalties to the loss function, reward-based objectives augment the loss with terms that encourage specific behaviors or properties. Pereyra et al. [29] introduced a confidence penalty to reduce overconfidence in predictive distributions, and Szegedy et al. [30] employed label smoothing to soften target distributions and mitigate excessive certainty. Müller et al. [31] systematically analyzed the effects of smoothing, summarizing its benefits for generalization and uncertainty calibration. Kim et al. [32] proposed negative learning, which leverages incorrect labels as training signals to improve robustness under noisy-label settings, while Ishida et al. [33] restructured objectives to utilize complementary labels instead of ground-truth labels.
@@
In sequence learning, several approaches sought to directly optimize rewards aligned with task-specific metrics. Ranzato et al. [34] proposed training with REINFORCE on sequence-level rewards, and Norouzi et al. [35] generalized maximum likelihood learning into reward-aligned training via reward-weighted distributions. In the domain of language model fine-tuning based on human preferences, Christiano et al. [36] and Ziegler et al. [37] demonstrated improving policies through reward models trained with human feedback, while Ouyang et al. [38] established their effectiveness at scale in instruction-following models. More recently, Rafailov et al. [39] proposed a supervised alternative that performs implicit reward maximization from preference data without explicit reinforcement learning.
@@
In representation learning, contrastive learning adopts reward structures that encourage similarity within positive pairs and separation across negative pairs. van den Oord et al. [40] proposed a contrastive loss that raises a lower bound on mutual information, thereby inducing representations with higher information content. Conversely, adversarial objectives designed to remove unwanted factors can be regarded as negative rewards. Ganin et al. [41] employed gradient reversal to invert domain-classification loss, preventing learned representations from encoding domain information, while Zhang et al. [42] introduced adversarial predictors of sensitive attributes to remove bias signals. In reinforcement learning, Haarnoja et al. [43] formulated the maximum entropy objective by adding entropy rewards to the reward function, thereby enhancing exploration and robustness.
2.3 Sample-size–dependent Regularization and Degrees-of-freedom Control
@@
The classical idea of adjusting the strength of regularization according to sample size and targeting the model’s effective degrees of freedom has been established in theories of information criteria, risk estimation, and cross-validation. Akaike [44] introduced the AIC, which penalizes model complexity (the effective number of parameters) with a constant term derived from bias correction of generalization error. Mallows [45] proposed the $C_{p}$ criterion as an unbiased estimator of expected prediction error. Stein [46] enabled data-adaptive selection of shrinkage strength through Stein’s Unbiased Risk Estimator (SURE), and Craven et al. [47] formalized generalized cross-validation (GCV) as a method for data-dependent selection of smoothing parameters. Wahba [48] systematized the equivalence between degrees of freedom and smoothness control in spline regularization, while Ye [49] defined generalized degrees of freedom (GDF) as the sum of sensitivities of fitted values to observations, showing that for linear smoothers this coincides with the trace of the hat matrix, thereby broadening the theoretical foundations of data-dependent complexity penalties. Hastie et al. [9] synthesized these criteria under a unified perspective and discussed operational rules in which penalty strength should decrease gradually with sample size to avoid underfitting and minimize prediction risk.
@@
In sparsity-inducing regularization, the interplay between degrees of freedom and penalty strength has also been analyzed in depth. Efron et al. [50] described the piecewise linear structure of Lasso solutions and their degrees-of-freedom dynamics via the LARS path, and Zou et al. [51] showed that under general conditions, the degrees of freedom of Lasso are well approximated by the number of active variables. Tibshirani et al. [52] rigorously characterized the degrees of freedom of Lasso-type problems under constraint and regularity assumptions, providing quantitative guidance for selecting regularization scales depending on sample size and signal strength. Donoho et al. [53] demonstrated that SURE-based thresholding leads to data-adaptive shrinkage levels that minimize risk, while Stone [54] provided theoretical support that cross-validation adaptively identifies the balance point between underfitting and overfitting as sample size increases.
@@
At the level of training procedures, early stopping has been shown to act as data-dependent regularization. Prechelt [55] empirically demonstrated that stopping rules based on validation error yield a regularization effect and suppress overfitting, while Yao et al. [56] established conditions under which early stopping in gradient descent is equivalent to Tikhonov regularization. Raskutti et al. [57] derived generalization bounds for nonparametric regression when the number of iterations is chosen in a sample-dependent manner, supporting the conclusion that more iterations or weaker explicit regularization become preferable as data size increases.
@@
In stochastic optimization, the noise scale induced by batch size and learning rate has been observed to act as an implicit regularization strength, and this effect must be scheduled with respect to sample size. Mandt et al. [58] approximated SGD as a stochastic differential equation, showing that the noise temperature mediates the trade-off between prior and posterior, effectively serving as a form of regularization. Smith et al. [59] explained the generalization optimum through scaling laws relating batch size, learning rate, and dataset size. Jastrzębski et al. [60] reported how combinations of dataset size, learning rate, and batch size influence flat-minimum exploration and generalization, while Keskar et al. [61] pointed out that large batch sizes can weaken implicit regularization and induce generalization gaps. Loshchilov et al. [62] decoupled weight decay from the optimization step to enhance schedulability, and Smith [63] proposed cyclical learning rates as a practical procedure for dynamically adjusting the strength of implicit regularization.
2.4 Safety in Classification and Training Stability